# The Milky Way as Seen by Classical Cepheids I: Distances Based on Mid-infrared Photometry

### v1

@@ 2.2 Mid-IR Photometry
where the $flux$ and $dflux$ are $W1$ flux and formal flux uncertainties as given in unWISE. The 0.02 in the second term defines a floor in the uncertainty of the unWISE $W1$ magnitudes due to unaccounted systematics. We arrive at this value using the differences between the unWISE and CatWISE $W1$ magnitudes for a large sample of sources, and estimating the additional uncertainty needed in unWISE to account for the observed standard deviation $(\sigma)$ of these differences for unWISE magnitudes $W1>8.5$ mag, on the assumption that the CatWISE uncertainties are correct. An additional Gaussian dispersion of 0.02 mag is needed to account for the observed $(\sigma)$ of the differences at magnitudes fainter than about $W1_{U}\simeq 10.5$ mag (see Figure 6), which we attribute to unWISE uncertainties. This gives perhaps a conservative estimate on the unWISE magnitude errors but, as we show in the following section, the uncertainties of our photometric distance moduli derived from the PL relation are still almost an order of magnitude larger than these estimated photometric errors.
@@
Figure 6: Standard deviation of the difference in $W1$ photometry between unWISE and catWISE shown in black, as a function of $W1$ from unWISE, for a field far away from the Galactic plane ( $205^{\circ}<l<225^{\circ}$ $10^{\circ}<b<20^{\circ}$) and $G<17$. The green curve is the nominal uncertainty in catWISE W1, while in blue we boost this value by adding in quadrature additional uncertainty $\sigma_{W1}$ from Equation 1, in order to scale up to the observed curve in black.
@@
Another source of uncertainty is related to estimating the mean brightness of Cepheids, taking into account their variability caused by pulsations. The unWISE and AllWISE source catalogs provide single brightness measurement performed on coadded frames, so no information about variability is contained in the reported photometric uncertainty. The mid-IR light curves for some Cepheids in our sample are available from the AllWISE Multiepoch Photometry Database, but they are often grouped around one or two epochs, which is equivalent to one or two light curve data points, given the fairly long pulsation periods of Cepheids. However, for some Cepheid light curves the number of epochs is sufficient to determine their mid-IR amplitudes and these have been used by Skowron et al. (2019a) to estimate the conversion factor between the mid-IR and optical amplitudes, as the latter are available for all Cepheids in the sample, contrary to mid-IR amplitudes. They found that the rms scatter in all mid-IR bands is half the rms scatter in the I-band and $\sim$30the rms scatter in the V-band (i.e., ${Amp}_{W1}=0.5\>{Amp}_{I}$ and ${Amp}_{W1}=0.3\>{Amp}_{V}$). We use the I-band amplitude values provided by the OGLE survey, which are reliable, as all have been verified by eye, to estimate $W1$ amplitude values. However, not all Cepheids in our sample have the optical amplitudes from OGLE, therefore for the ones that do not, we use peak_to_peak_g amplitudes from the gaiadr3.vari_cepheid table. We found that the conversion factor between Gaia BP band and WISE $W1$ is $0.32$ such that ${Amp}_{\rm W1}=0.32\>{Amp}_{\rm BP}$.
@@
The mid-IR light curves of Cepheids are typically sinusoidal. The uncertainty of estimating the mean value for a sinusoidal from one random epoch is equal to $0.7$ of its amplitude (i.e. $Amp_{sin}/\sqrt{2}$). This means that if we estimate the mean brightness of a Cepheid from one random epoch, its uncertainty is $Amp_{W1}/(2\sqrt{2})\simeq 0.35\>Amp_{W1}$, since the amplitude of the sinusoidal is half that of the peak-to-peak amplitude of a Cepheid. However, magnitudes from both AllWISE and unWISE Catalogs are based on co-added images taken over multiple epochs. AllWISE was produced from two full-sky scans [^6] AllWISE release notes, while the observational baseline of unWISE was about five times longer (Schlafly et al., 2019). Simulations show, that having two points on a sinusoidal reduces the uncertainty of estimating the mean to $0.5$ of the amplitude, while having ten points to $0.22$ of the amplitude. As a result, the uncertainty in estimating the mean brightness of a Cepheid from resulting from pulsations will be $0.25\>Amp_{W1}$ for AllWISE and $0.11\>Amp_{W1}$ for unWISE.
@@ 3.1.4 Final distances and their uncertainties
Figure 8: Photometric distances derived from the $W1$ magnitudes for 2780 Cepheids for which both AllWISE and unWISE photometry was available. Orange represents bright Cepheids with $W1_{U}<8.25$ mag, while the remaining are shown in blue.
As shown in Figure 8, after the magnitude offsets from Eq. 3.1 have been applied, both AllWISE and unWISE distances seem to be consistent with each other, even at bright magnitudes (see also Figure 5). The median differences between AllWISE and unWISE distances are very small, i.e. $-0.002$ kpc for the whole sample, $-0.01$ kpc for Cepheids brighter than $W1_{U}=8.25$ mag and $0.001$ kpc for fainter than $W1_{U}=8.25$ mag. The respective standard deviations are $0.472$, $0.205$, and $521$ kpc. We have therefore decided to also derive $W1$-based distance moduli using both the AllWISE and unWISE magnitudes, by averaging the magnitudes when both are available. The uncertainty of the averaged WISE magnitudes was calculated as half the sum in quadrature of $\sigma_{W1}$ from AllWISE and unWISE. The final photometric uncertainty was calculated with Eq. 2, using the same amplitude scaling factor as for unWISE data. We then calculated the distance based on the averaged magnitudes iteratively, as described above.
The final count of Cepheids with $W1$-based distances is 3425, of which 2780 have both AllWISE and unWISE $W1$ magnitudes, while 277 have only AllWISE and 368 have only unWISE $W1$ photometry. In other words, for 2780 Cepheids three different distance estimates are available, while for 645 stars only one distance value based on either AllWISE or unWISE data.
The full set of distances and the distance moduli are available online777link here, while in Table 1 we describe the columns in our catalog.
@@
No. Column Description Unit
1 survey_source Source catalog as reported by P21 -
2 survey_source_id Source ID in the Gaia survey_source catalog -
3 gaia_id Gaia source ID -
4 phot_g_mean_mag Gaia $G$ magnitude magnitude
5 glon galactic longitude degrees
6 glat galactic latitude degrees
7 mode Mode of variability, from P21 -
8 period Period of variability, from P21 days
9 w1_all $W1$ magnitude from AllWISE, after corrections magnitude
10 mu_all distance modulus based on AllWISE magnitude
11 dmu_all uncertainty in distance modulus based on AllWISE magnitude
12 d_all distance based on AllWISE parsec
13 dd_all uncertainty in distance based on AllWISE parsec
14 a_all extinction in $W1$ from AllWISE magnitude
15 w1_un $W1$ magnitude from unWISE magnitude
16 mu_un distance modulus based on unWISE magnitude
17 dmu_un uncertainty in distance modulus based on UnWISE magnitude
18 d_un distance based on UnWISE parsec
19 dd_un uncertainty in distance based on UnWISE parsec
20 a_un extinction in $W1$ from unWISE magnitude
21 w1_av $W1$ magnitude averaged over AllWISE & unWISE magnitude
22 mu_av distance modulus based on averaged W1
23 dmu_av uncertainty in distance modulus based on averaged W1 magnitude
24 d_av distance based on averaged W1 parsec
25 dd_av uncertainty in distance based on averaged W1 parsec
26 a_av extinction in $W1$ based on averaged W1 magnitude
27 age Age estimate Myr
28 Q quality indicator -
@@ 3.2 Validation of photometric distances
where $\varpi$ is the parallax in $\mu\rm{as}$. Figure 9 shows the difference in the distance moduli $\Delta\mu=\mu_{\varpi}-\mu_{W1}$ as a function of the unWISE $W1$ magnitudes. While the majority of the sources show overall consistent distances, there is a conspicuous set of sources whose $\Delta\mu<2$ mag. In particular, we find 20 stars classified as Cepheids according to their light curves, but whose astrometric distances show they cannot be Cepheids. All of these stars are classified as first-overtone oscillators. We inspected the light curve of these sources and found that, after including more recent photometric observations, a number now appear to be spotted stars, while others cannot be clearly classified. However, at least five still appear to be Cepheids (P. Pietrukowicz, private communication).
@@
where $\varpi_{0}$ is the parallax zeropoint (i.e. offset), and $\epsilon_{\varpi}$ is a random normal deviate with a standard deviation that is the sum in quadrature of the associated $\sigma_{\varpi}$, and an additional $\sigma_{0}$ to account for the variation of the parallax offset for our subsample (Lindegren et al., 2021b). That is, we assume that the parallax offset has a mean value $\varpi_{0}$, and varies from this as a Gaussian with a standard deviation of $\sigma_{0}$. In short, the set of 910 resampled parallaxes $\varpi$ is sampled from the distribution $\mathcal{N}(\varpi\,|\,\varpi_{true}-\varpi_{0},\sigma^{2}_{\varpi}+\sigma^{2}_{0})$ where $\mathcal{N}(x\,|\,\langle x\rangle,\sigma^{2})$ is a normal distribution with mean $\langle x\rangle$ and standard deviation $\sigma$. Our resulting distribution of resampled parallaxes will of course be smoother than the actual data since we are taking $\varpi_{true}$ from a bootstrap of the observed parallaxes, but we are not trying to model or replicate the distribution of parallaxes. Our procedure is a simple way to generate a ”realistic” distribution of model observations with (conveniently) similar correlations as is in the actual data, without needing to explicitly model these correlations and the parallax error distribution. For example, more distant (fainter) sources will have larger errors in our bootstrap sample just as in the real data since we drew $\{\varpi,\sigma_{\varpi}\}$ pairs from the data itself.
@@
To find a more refined estimate of our parameters that does not rely on binning the data, we instead maximize the Likelihood of our dataset $\{\Delta\mu_{i}\}$ using Markov chain Monte Carlo (MCMC), as implemented in the emcee package (Foreman-Mackey et al., 2013). We again generate a model distribution of $\Delta\mu$ from a single set of $\{\varpi_{true},\sigma_{\varpi}\}$ bootstrapped from our data set, but oversampling the noise by a factor of 20. From this distribution we derive a probability distribution function (PDF) using a normalized kernel density estimator from the sklearn package Pedregosa et al. (2011), using a Gaussian kernel with a bandwidth of 0.05 mag and sampled with 1000 points in $\Delta\mu$ over the full range of our stochastic sample generated from our bootstrap sample, typically covering the range [-1,2]. The probability of each data point $\Delta\mu_{i}$ is found from the PDF by linear interpolation. We then find the log likelihood ($\log L$) of our dataset $\{\Delta\mu_{i}\}$ by summing the log probabilities of the individual points. As our dataset does include outliers that fall outside the range of our stochastically modeled PDF but contribute to $\log L$, we found it necessary to ”stabilize” $\log L$ by setting a minimum value for the probability to $10^{-8}$. We estimate our parameters by using MCMC with 20000 steps and 24 walkers, and using the $\chi^{2}$ estimated values above as a first guess. In addition we found it necessary to limit the range of possible parameter values explored by the MCMC by setting the $logL$ to negative infinity if the explored value of any of the three parameters falls outside the prescribed range, reported in the last column of Table 2, which summarizes the resulting best estimates and their uncertainties from the MCMC samples after removing the first 4000 steps to allow for burn in. Figure 10 shows the resulting corner plot of possible parameter values and Figure 11 shows the observed $\Delta\mu$ distribution and its modeled PDF using these final parameter estimates from the MCMC maximizing the Likelihood.
@@
Table 2: Estimated parameters
@@
We note that the offset of the mode of the $\Delta\mu$ distribution from zero, as well as the asymmetry of the distribution, can only be attributed to the astrometric distance moduli errors, the offset being primarily due to the parallax offset $\varpi_{0}$ and the asymmetry due the parallax errors. Our estimated parallax offset is comparable with other estimates in the literature, as discussed below in Section 4. To summarize, we estimate the uncertainty of our photometric distance moduli $\sigma_{\mu}$ to be $0.13\pm 0.04$ mag, that is, to have relative distance errors of less than 13%. Meanwhile the mean $\sigma_{\mu}$ of the subsample of stars used is 0.09 mag. Our verification using the measured parallaxes thus suggest that our estimated errors on the distance moduli are somewhat under-estimated, at least for this subset of the nearest Cepheids. This under-estimation cannot only in part be attributed to neglecting the extinction uncertainties, as the mean $A_{W1}$ of this subsample is only 0.12 mag so that even a 50% uncertainty in the extinction would only contribute an additional 0.02 mag to the distance modulus uncertainties. However, it is worth noting from Figure 10 that $\sigma_{0}$ and $\sigma_{\mu}$ are correlated and, more fundamentally, our analysis is based on the assumption that there is no systematic error in the magnitudes.
Now that we have characterized our photometric distance modulus errors, we can provide a quantification of the consistency of the astrometric and photometric distances, or rather, the consistency of the astrometric and photometric parallaxes, where the photometric parallax (in $\mu\rm{as}$) is
@@
We find that $Q<5$ removes the 20 anomalous sources with good parallaxes and inconsistent astrometric/photometric distance moduli ($\Delta\mu<-2$mag) mentioned above. Applying this criteria to the complete sample of Cepheids removes an additional 43 sources, that is 63 out of 3425 (1.8%) of our sample with distance moduli. In Figure 12 we show the cumulative distribution of $Q$, which can be used as a quality indicator for selecting subsamples.
Figure 12: Cumulative distribution of the normalized difference in the $W1$-based photometric and parallax-based astrometric parallaxes ($|\Delta\varpi/\sigma_{\Delta\varpi}|$).
To test whether our uncertainties are indeed underestimated due to neglecting extinction uncertainties, we checked whether $Q$ was correlated with extinction $A_{W1}$. As the uncertainty in extinction should be proportional to extinction, higher $Q$ values should result from systematically underestimating the uncertainties with higher extinction. However we found no evidence of a correlation between $Q$ and $A_{W1}$.

### v2

@@ 2.2.2 Photometric uncertainties
where the $flux$ and $dflux$ are $W1$ flux and formal flux uncertainties as given in unWISE. The 0.02 in the second term defines a floor in the uncertainty of the unWISE $W1$ magnitudes due to unaccounted systematics. We arrive at this value using the differences between the unWISE and CatWISE $W1$ magnitudes for a large sample of sources, and estimating the additional uncertainty needed in unWISE to account for the observed standard deviation $(\sigma)$ of these differences for unWISE magnitudes $W1>8.5$ mag, on the assumption that the CatWISE uncertainties are correct. An additional Gaussian dispersion of 0.02 mag is needed to account for the observed $\sigma$ of the differences at magnitudes fainter than about $W1_{U}\simeq 10.5$ mag (see Figure 7), which we attribute to unWISE uncertainties. This gives perhaps a conservative estimate on the unWISE magnitude errors but, as we show in the following section, the uncertainties of our photometric distance moduli derived from the PL relation are still almost an order of magnitude larger than these estimated photometric errors.
@@
Figure 7: Standard deviation of the difference in $W1$ photometry between unWISE and catWISE shown in black, as a function of $W1$ from unWISE, for a field far away from the Galactic plane ( $205^{\circ}<l<225^{\circ}$ $10^{\circ}<b<20^{\circ}$) and $G<17$ mag. The green curve is the nominal uncertainty in catWISE $W1$, while in blue we boost this value by adding in quadrature additional uncertainty $\sigma_{W1}$ from Eq. 1, in order to scale up to the observed curve in black.
@@
Another source of uncertainty arises when estimating the mean brightness of Cepheids, due to their variability. That is, determining the mean brightness from a few random measurements taken over time introduces a larger uncertainty for a variable star than for a constant one, and the greater the amplitude of variability, the greater the uncertainty. The unWISE and AllWISE source catalogs provide single brightness measurements performed on coadded frames, so no information about the amplitude of variability is contained in the reported photometric uncertainty. The mid-IR light curves for some Cepheids in our sample are available from the AllWISE Multiepoch Photometry Database, but they are often grouped around one or two epochs, which is equivalent to one or two light curve data points, given the fairly long pulsation periods of Cepheids. However, for some Cepheid light curves the number of epochs is sufficient to determine their mid-IR amplitudes and these have been used by Skowron et al. (2019a) to estimate the conversion factor between the mid-IR and optical amplitudes, as the latter are available for all Cepheids in the sample, contrary to mid-IR amplitudes. They found that the rms scatter in all mid-IR bands is half the rms scatter in the I-band and $\sim$30the rms scatter in the V-band (i.e., ${Amp}_{W1}=0.5\>{Amp}_{I}$ and ${Amp}_{W1}=0.3\>{Amp}_{V}$). We use the I-band amplitude values provided by the OGLE survey, which are reliable, as all have been verified by eye, to estimate $W1$ amplitude values. However, not all Cepheids in our sample have the optical amplitudes from OGLE, therefore for the ones that do not, we use peak_to_peak_bp amplitudes from the gaiadr3.vari_cepheid table. We found that the conversion factor between Gaia BP band and WISE $W1$ is $0.32$ such that ${Amp}_{W1}=0.32\>{Amp}_{BP}$.
@@
The mid-IR light curves of Cepheids are typically sinusoidal. The uncertainty of estimating the mean value for a sinusoidal from one random epoch is equal to $0.7$ of its amplitude (i.e. $Amp_{sin}/\sqrt{2}$). This means that if we estimate the mean brightness of a Cepheid from one random epoch, its uncertainty is $Amp_{W1}/(2\sqrt{2})\simeq 0.35\>Amp_{W1}$, since the amplitude of the sinusoidal is half that of the peak-to-peak amplitude of a Cepheid. However, magnitudes from both AllWISE and unWISE Catalogs are based on co-added images taken over multiple epochs. AllWISE was produced from two full-sky scans [^7] AllWISE release notes, while the observational baseline of unWISE was about five times longer (Schlafly et al., 2019). Simulations show, that having two points on a sinusoidal reduces the uncertainty of estimating the mean to $0.5$ of the amplitude, while having ten points to $0.22$ of the amplitude. As a result, the uncertainty in estimating the mean brightness of a Cepheid resulting from pulsations will be $0.25\>Amp_{W1}$ for AllWISE and $0.11\>Amp_{W1}$ for unWISE.
@@ 3.1.3 Extinction correction
Figure 9: Extinction in $W1$ as a function of $G-W1$ color, shown for the Cepheids from this study as black ($-90^{\circ}<l<90^{\circ}$) and blue ($90^{\circ}<l<-90^{\circ}$) points. Orange line shows the prediction for $R_{V}=3.1$ assuming the extinction curve from Gordon et al. (2023), where we adopt a Cepheid intrinsic mean color of $(G-W1)_{0}=1.5$ mag. Cyan and red lines show the predictions for $R_{V}=2.6$ and $R_{V}=3.6$, respectively.
We note that $A_{W1}/A_{K_{S}}$ does not significantly vary over the effective temperature range of Classical Cepheids, but varies between 0.41 to 0.50 for $R_{V}$ from 2.6 to 3.6, that is at most 20% over this entire range, or 10% from the mean value at $R_{V}=3.1$. Therefore, the spread in $A_{W1}$ for a given $G-W1$ color in Figure 9 is primarily due to the variation of $R_{V}$, as shown with cyan and red lines representing $R_{V}=2.6$ and $R_{V}=3.6$, respectively. Additionally, part of the spread in Figure 9 can be attributed to the width of the instability strip. Notably, even when considering a much broader range of $R_{V}$ from 2.6 to 5.5, we arrive at $A_{W1}/A_{K_{S}}$ between 0.41 to 0.53, corresponding to a variation of about 13% from its mean value at $R_{V}=3.1$.
The mean extinction $A_{W1}$ for our sample is 0.22 mag, while 90% stars have $A_{W1}\leq 0.45$ mag, so that we expect an unaccounted variation of $R_{V}$ to introduce an uncertainty of $<0.045$ mag (typically $0.02-0.03$ mag) in our estimated extinctions, which would translate to $<2.1\%$ relative distance uncertainty (typically $1\%$).
In addition to the uncertainty in the estimated extinction due to an unknown variation in $R_{V}$, there is also the uncertainty coming from the extinction maps. Unfortunately $mwdust$ does not provide an extinction error for specific coordinates. We find that the majority of the extinction from $mwdust$ for our Cepheids come from either Marshall et al. (2006) or Green et al. (2019). The typical $A_{W1}$ uncertainties that come from the Marshall map are about 10%, with the larger uncertainties being from the brighter nearby sources with a small amount of extinction. However, the typical relative uncertainties cited in the Green map are 3.5%, which seem overly optimistic, therefore we assume a relative $A_{W1}$ uncertainty of 10% for both maps. The extinction error from a possible variation of $R_{V}$ between 2.6 and 3.6 of 10%, combined in quadrature with this 10% uncertainty, gives a total possible mean $A_{W1}$ uncertainty of 14%, which results in 1.4% maximum relative distance uncertainty for $A_{W1}=0.22$ mag and 2.9% for $A_{W1}=0.45$ mag.
@@ 3.1.4 Final distances and their uncertainties
Figure 10: Photometric distances derived from the $W1$ magnitudes for 2780 Cepheids for which both AllWISE and unWISE photometry was available. Orange represents bright Cepheids with $W1_{U}<8.25$ mag, while the remaining ones are shown in blue.
As shown in Figure 10, after the magnitude offsets from Eq. 3.1 have been applied, both AllWISE and unWISE distances seem to be consistent with each other, even at bright magnitudes (see also Figure 5). The median differences between AllWISE and unWISE distances are very small, i.e. $-0.002$ kpc for the whole sample, $-0.01$ kpc for Cepheids brighter than $W1_{U}=8.25$ mag and $0.001$ kpc for fainter than $W1_{U}=8.25$ mag. The respective standard deviations are $0.472$, $0.205$, and $0.521$ kpc. We have therefore decided to also derive $W1$-based distance moduli using both the AllWISE and unWISE magnitudes, by averaging the magnitudes when both are available. The uncertainty of the averaged WISE magnitudes was calculated as half the sum in quadrature of $\sigma_{W1}$ from AllWISE and unWISE. The final photometric uncertainty was calculated with Eq. 2, using the same amplitude scaling factor as for unWISE data. We then calculated the distance based on the averaged magnitudes iteratively, as described above.
The final count of Cepheids with $W1$-based distances is 3424, of which 2780 have both AllWISE and unWISE $W1$ magnitudes, while 277 have only AllWISE and 368 have only unWISE $W1$ photometry. In other words, for 2780 Cepheids three different distance estimates are available, while for 645 stars only one distance value based on either AllWISE or unWISE data.
The full set of distances and the distance moduli are available online,while in Table 1 we describe the columns in our catalog.
@@
No. Column Description Unit Example
1 survey_source Source catalog as reported by P21 - GCVS
2 survey_source_id Source ID in the survey_source catalog - CG____Cas
3 gaia_id Gaia source ID - 429385923752386944
4 glon Galactic longitude degrees 116.845566
5 glat Galactic latitude degrees -1.315165
6 mode Mode of variability, from P21 - F
7 period Fundamentalized Period of variability, from P21 days 4.36540040
8 w1_all $W1$ magnitude from AllWISE, after corrections magnitude 7.961
9 mu_all distance modulus based on AllWISE magnitude 12.494
10 dmu_all uncertainty in distance modulus based on AllWISE magnitude 0.112
11 d_all distance based on AllWISE parsec 3154
12 dd_all uncertainty in distance based on AllWISE parsec 162
13 a_all extinction in $W1$ from AllWISE magnitude 0.079
14 w1_un $W1$ magnitude from unWISE magnitude 7.989
15 mu_un distance modulus based on unWISE magnitude 12.522
16 dmu_un uncertainty in distance modulus based on unWISE magnitude 0.090
17 d_un distance based on unWISE parsec 3194
18 dd_un uncertainty in distance based on unWISE parsec 133
19 a_un extinction in $W1$ from unWISE magnitude 0.079
20 w1_av $W1$ magnitude averaged over AllWISE & unWISE magnitude 7.975
21 mu_av distance modulus based on averaged $W1$ magnitude 12.508
22 dmu_av uncertainty in distance modulus based on averaged $W1$ magnitude 0.089
23 d_av distance based on averaged $W1$ parsec 3174
24 dd_av uncertainty in distance based on averaged $W1$ parsec 131
25 a_av extinction in $W1$ based on averaged $W1$ magnitude 0.079
26 age Age estimate Myr 128
27 Q quality indicator - 1.528
@@ 3.2 Validation of photometric distances
where $\varpi$ is the parallax in mas. Figure 11 shows the difference in the distance moduli $\Delta\mu=\mu_{\varpi}-\mu_{W1}$ as a function of the unWISE $W1$ magnitudes. While the majority of the sources show overall consistent distances, there is a conspicuous set of sources whose $\Delta\mu<2$ mag. In particular, we find 20 stars classified as Cepheids according to their light curves, but whose astrometric distances show they cannot be Cepheids. All of these stars are classified as first-overtone oscillators. We inspected the light curve of these sources and found that, after including more recent photometric observations, a number now appear to be spotted stars, while others cannot be clearly classified. However, at least five still appear to be Cepheids, although this classification may change as more data are available (P. Pietrukowicz, private communication).
@@
where $\varpi_{0}$ is the parallax zeropoint (i.e. offset), and $\epsilon_{\varpi}$ is a random normal deviate with a standard deviation that is the sum in quadrature of the associated $\sigma_{\varpi}$, and an additional $\sigma_{0}$ to account for the variation of the parallax offset for our subsample (Lindegren et al., 2021b, hereafter L21). That is, we assume that the parallax offset has a mean value $\varpi_{0}$, and varies from this as a Gaussian with a standard deviation of $\sigma_{0}$. In short, the set of 910 resampled parallaxes $\varpi$ is sampled from the distribution $\mathcal{N}(\varpi\,|\,\varpi_{true}-\varpi_{0},\sigma^{2}_{\varpi}+\sigma^{2}_{0})$ where $\mathcal{N}(x\,|\,\langle x\rangle,\sigma^{2})$ is a normal distribution with mean $\langle x\rangle$ and standard deviation $\sigma$. Our resulting distribution of resampled parallaxes will of course be smoother than the actual data since we are taking $\varpi_{true}$ from a bootstrap of the observed parallaxes, but we are not trying to model or replicate the distribution of parallaxes. Our procedure is a simple way to generate a ”realistic” distribution of model observations with (conveniently) similar correlations as is in the actual data, without needing to explicitly model these correlations and the parallax error distribution. For example, more distant (fainter) sources will have larger errors in our bootstrap sample just as in the real data since we drew $\{\varpi,\sigma_{\varpi}\}$ pairs from the data itself.
@@
To find a more refined estimate of our parameters that does not rely on binning the data, we instead maximize the Likelihood of our dataset $\{\Delta\mu_{i}\}$ using Markov chain Monte Carlo (MCMC), as implemented in the emcee package (Foreman-Mackey et al., 2013). We again generate a model distribution of $\Delta\mu$ from a single set of $\{\varpi_{true},\sigma_{\varpi}\}$ bootstrapped from our data set, but oversampling the noise by a factor of 20. From this distribution we derive a probability distribution function (PDF) using a normalized kernel density estimator from the sklearn package Pedregosa et al. (2011), using a Gaussian kernel with a bandwidth of 0.05 mag and sampled with 1000 points in $\Delta\mu$ over the full range of our stochastic sample generated from our bootstrap sample, typically covering the range [-1,2]. The probability of each data point $\Delta\mu_{i}$ is found from the PDF by linear interpolation. We then find the log likelihood ($\log L$) of our dataset $\{\Delta\mu_{i}\}$ by summing the log probabilities of the individual points. As our dataset does include outliers that fall outside the range of our stochastically modeled PDF but contribute to $\log L$, we found it necessary to ”stabilize” $\log L$ by setting a minimum value for the probability to $10^{-6}$. We estimate our parameters by using MCMC with 20000 steps and 24 walkers, and using the $\chi^{2}$ estimated values above as a first guess. In addition we found it necessary to limit the range of possible parameter values explored by the MCMC by setting the $logL$ to negative infinity if the explored value of any of the three parameters falls outside the prescribed range. These ranges are reported in the last column of Table 2, which summarizes the resulting best estimates and their uncertainties from the MCMC samples after removing the first 4000 steps to allow for burn in. Figure 12 shows the resulting corner plot of possible parameter values and Figure 13 shows the observed $\Delta\mu$ distribution and its modeled PDF using these final parameter estimates from the MCMC maximizing the Likelihood.
@@
Table 2: Estimated parameters for parallax zeropoint, parallax zeropoint variance and photometric uncertainties.
@@
We note that the offset of the mode of the $\Delta\mu$ distribution from zero, as well as the asymmetry of the distribution, can only be attributed to the astrometric distance moduli errors, the offset being primarily due to the parallax offset $\varpi_{0}$ and the asymmetry due the parallax errors. Our estimated parallax offset is comparable with other estimates in the literature, as discussed below in Section 4. To summarize, we estimate the uncertainty of our photometric distance moduli $\sigma_{\mu}$ to be $0.13\pm 0.04$ mag, that is, to have relative distance errors of less than 6% in the magnitude range considered. Meanwhile the mean uncertainty of the distance moduli of this subsample of stars, derived in the previous section, is 0.09 mag. Our verification using the measured parallaxes thus suggest that our estimated errors on the distance moduli are somewhat under-estimated, at least for this subset of the nearest Cepheids. This under-estimation of the uncertainties cannot be attributed solely to the neglected extinction uncertainties, as the mean $A_{W1}$ of this subsample is only 0.12 mag so that even a 50% uncertainty in the extinction would only contribute an additional 0.02 mag to the distance modulus uncertainties. However, it is worth noting from Figure 12 that $\sigma_{0}$ and $\sigma_{\mu}$ are correlated and, more fundamentally, our analysis is based on the assumption that there is no systematic error in the magnitudes.
Now that we have characterized our photometric distance modulus errors, we can provide a quantification of the consistency of the astrometric and photometric distances, or rather, the consistency of the astrometric and photometric parallaxes, where the photometric parallax (in mas) is
@@
We find that $Q<5$ removes the 20 anomalous sources with good parallaxes and inconsistent astrometric/photometric distance moduli ($\Delta\mu<-2$ mag) mentioned above. Applying this criteria to the complete sample of Cepheids removes an additional 43 sources, that is 63 out of 3424 (1.8%) of our sample with distance moduli. In Figure 14 we show the cumulative distribution of $Q$, which can be used as a quality indicator for selecting subsamples.
Figure 14: Cumulative distribution of the normalized difference in the $W1$-based photometric and parallax-based astrometric parallaxes ($Q=|\Delta\varpi/\sigma_{\Delta\varpi}|$).
To test whether our uncertainties are underestimated due to neglecting extinction uncertainties, we also checked whether $Q$ was correlated with extinction $A_{W1}$. As the uncertainty in extinction should be proportional to extinction, higher $Q$ values should result from systematically underestimating the uncertainties with higher extinction. However we found no evidence of such a correlation between $Q$ and $A_{W1}$.